consider a room of 15m x 15m x 5 m. If the pressure in the room is 101 kPa and the temperature is 35°C, what is the mass of the air in the room?
Answer #1

Given data ,

Pressure , P = 101 KPa

Temperature , T = 35 C = 35 +273 = 308 K

Volume , V = 15*15*5 = 1125 m^3

We know from gas equation ,

PV = m R T

m = PV/ RT

m= (101 * 1125 ) / (0.287 * 308 ) [ for air , gas conscent ,R = 0.287 KJ/Kg K ]

m = 1285.40 Kg

Hence mass of the air in the room is 1285.40 Kg